OpenStudy (anonymous):

A little help? Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

it costs them $0.50 to buy each flower they then turn around and sell each flower at $0.50 + 40% of $0.50 = 0.50 + 0.4*0.5 = 0.70 dollars

OpenStudy (anonymous):

you must buy and sell 2500 flowers to get a profit of $500

OpenStudy (anonymous):

you want the profit to be $500, which means P = 0.20x 500 = 0.20x ... plug in P = 500 then solve for x

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome